OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. m = 7, point (-2, 0)

OpenStudy (johnweldon1993):

So you may use slope intercept form here y - y1 = m(x - x1) You have m = slope = 7 and you have a point (x,y)(-2,0) so plug in your 'y' for y1 and plug in your 'x' for x1 and plug in your 'm' for m what do you get?

OpenStudy (anonymous):

6

OpenStudy (johnweldon1993):

not quite...I literally meant plug them into that equation y - y1 = m(x - x1) so y - 0 = 7(x + 2) and simplify that...what do you get now?

OpenStudy (anonymous):

-2 - 0?

OpenStudy (anonymous):

ugh i dont get this

OpenStudy (johnweldon1993):

don't get aggravated you'll get it :) have you worked with variables before? meaning x's and y's?

OpenStudy (anonymous):

yes, i get that you plug in the numbers, but how exactly?

OpenStudy (johnweldon1993):

It's important to realize...that for an equation of a line...you are going to have something that looks like *in this case* y = mx + b So when you start off with y - y1 = m(x - x1) you need to make this equation look like that equation so when you plug in 7 for m -2 for x1 and 0 for y1 you get y - 0 = 7(x + 2) now you need to make this look like y = mx +b so we should distribute that 7 into the parenthesis y = 7x + 14 And that will be your equation of the line you are looking for! If anything is unclear let me know and I'll break it down more :)